OpenStudy (anonymous):

Please help :/ Find f^-1(x) if f(x)=5x-2/4 f^-1(x)= ?

OpenStudy (anonymous):

If f(x) = −4x − 2, find f(−3) f(-3) = -4*(-3) - 2 = 12 - 2 = 10 If h(x) = 2x + 1, find the value of x if h(x) = 11. 2x + 1 = 11 2x = 10 x = 5 If f(x) = x + 4, find f(3). f(3) = 3 + 4 = 7 If g(x) = 2(x − 4), find the value of x if g(x) = 20. 2(x − 4) = 20 x - 4 = 10 x = 14 hope this helps u

OpenStudy (anonymous):

:/ Thank you so much but that doesn't seem to fit my options: a. 4x-5 ---- 2 b. 4x+2/5 c. 4x-2/5 d. 4x+5/2

OpenStudy (anonymous):

Rohangrr, you there?

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Thanks for what?

OpenStudy (anonymous):

for ur comment!! :D have a nice day

OpenStudy (anonymous):

You're not making sense...

OpenStudy (anonymous):

5x-2/4 or (5x-4)/2

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

x/5+0.1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I told you the answer

OpenStudy (anonymous):

:? But that doesn't match a., b., c., or d.